Solve
5/4x = 3/x - 2

Answer by ptaylor(2198)

5/4x = 3/x - 2 multiply each term by 4x (the LCM)
5=12-8x subtract 12 from each side
5-12=12-12-8x collect like terms
-7=-8x divide both sides by -8
x=7/8----------------------ans
CK
5/4(7/8)=3/(7/8)-2 or
10/7=24/7-2 subtract 24/7 from each side
10/7-24/7=-2
-14/7=-2
-2=-2
Hope this helps---ptaylor